# Ryan Zarick     Bryan Pellegrino     Isaac Zhang     Thomas Kim     Caleb Banister LayerZero Labs

### v2

@@ Abstract
footnotetext: Copyright © 2023 LayerZero Labs Ltd. All rights reserved.footnotetext: (2024-01-23) Version 1.1: Fix typos.
@@ 3.0.1 LayerZero packet transmission
In step \⃝raisebox{-0.9pt}{1}, the sender calls lzSend on the source chain LayerZero Endpoint, specifying the message payload and the path. This path is associated with an independent censorship resistant channel, and is composed of the sender application address, the source Endpoint ID, the recipient application address, and the destination Endpoint ID.
@@
In step \⃝raisebox{-0.9pt}{2}, the configured DVNs each independently verify the packet on the destination MessageLib; for ULN, this constitutes storing the hash of the packet payload. After a threshold of DVNs verify the payload (see Section 3.2.1), a worker (e.g., executor, DVN, user) commits the packet to the Endpoint in step \⃝raisebox{-0.9pt}{3}. The Endpoint checks that the payload verification reflects the OApp-configured Security Stack before committing to the lossless channel.
Finally, in step \⃝raisebox{-0.9pt}{4}, an Executor calls lzReceive on the committed message to execute the Receiver OApp logic on the packet. Step \⃝raisebox{-0.9pt}{4} will revert to prevent censorship if the channel cannot guarantee lossless exactly-once delivery.
@@ 3.1 LayerZero Endpoint
nilify Called by the Executor to receive a message from the channel.
@@
burn Called by the Executor to receive a message from the channel.
@@ 3.2.1 Ultra Light Node
[…]

### v3

@@ Abstract
footnotetext: Copyright © 2023 LayerZero Labs Ltd. All rights reserved.footnotetext: (2024-01-23) Version 1.1: Fix typos.footnotetext: (2024-03-07) Version 1.2: Fix figure typo.footnotetext: (2025-05-13) Version 1.3: Fix typo in Table 3
@@ 3.0.1 LayerZero packet transmission
[…]
@@
[…]
@@ 3.1 LayerZero Endpoint
nilify Called by the receiver to temporarily invalidate a nonce.
@@
burn Called by the receiver to delete and skip a nonce.
@@ 3.2.1 Ultra Light Node
We illustrate an example of the Ultra Light Node verification semantics in Figure 6. The OApp Security Stack includes required DVN (DVNA) and $N-1$ optional DVNs (DVNB, DVNC, …) with an OptionalThreshold of 1. This gives DVNA “veto” power, and also requires at least one of the optional DVNs to verify the packet before it can be committed. Nonce 1 has been committed to the messaging channel (Verified). Nonces 2, 3, and 6 are committable (verifiable) as the Security Stack has been fulfilled, but are not committed until an executor calls commitVerification. Nonces 4 and 5 are not committable because they have not fulfilled the required verifier set and OptionalThreshold respectively.